Model Theory in its First Century

Wilfrid Hodges

Model theory is a set of theorems, and of people. But more fundamentally it is a collection of ways of looking at mathematical structures. It gained its name in a paper of Alfred Tarski in 1954, which followed on from addresses given by Tarski and Abraham Robinson to the 1950 International Congress of Mathematicians. Several strands of work from the first half of the century came together in those addresses, and in the second half of the century many lines of research traced back to them.

Address: School of Mathematical Sciences, Queen Mary and Westfield College, Mile End Road, London E1 4NS, England

E-mail: w.hodges@qmw.ac.uk